OpenStudy (anonymous):

If log (with base 5)x=8, what is the value of x?

OpenStudy (cwrw238):

x = 5^8

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm log_ab=c }\) \(\Huge\color{lime}\rightarrow\) \(\large\color{midnightblue}{ \rm a^c=b }\)

OpenStudy (anonymous):

remember this: \[\log _{a} x = y \Longrightarrow x = a^y\]

OpenStudy (anonymous):

oh @SolomonZelman said this at the same time ;)

OpenStudy (anonymous):

it says here though that the answers are a) 1.29 b)1.52 c)32,768 d)390,625

OpenStudy (anonymous):

okay nevermind I got d

jimthompson5910 (jim_thompson5910):

As the others are saying, you convert to exponential form to get x = 5^8 Then notice how the powers of 5 are 5^1 = 5 5^2 = 25 5^3 = 125 5^4 = 625 Each power of 5 ends with 5 So 5^8 will end with 5